# Auxiliary Discrminator Sequence Generative Adversarial Networks (ADSeqGAN) for Few Sample Molecule Generation

### v1

@@ 4.1 Nucleic Acid Binder Generation
To distinguish nucleic acid binders from protein-targeting small molecules, we computed molecular descriptors for a dataset of compounds and evaluated their discriminatory power using logistic regression. Following an initial regression-based analysis, we manually selected 18 descriptors with distinct physicochemical properties that showed relevance in distinguishing these two classes. A random forest classifier trained on these selected descriptors achieved an AUC of 0.91 (Figure S1), indicating strong classification performance.
@@
Figure 2: ADSeqGAN training results on the nucleic acid and protein dataset. $A.$ The yield of nucleic acid binders using "NA" as the input label changes with the increase of epoch. Yield is calculated by $unique\_ratio\times verified\_ratio\times NA\_ratio$. $B.$ NLDock docking results of generated molecules with various nucleic acid targets. Use flexible and local modes. 1C9Z and 1EEL are DNA targets, while 2L94 and 6XB7 are RNA targets. The grey balls are original binders with native conformations, and the red ones are generated samples. $C.$ Plot of label responsiveness with the increase of epoch. The responsiveness metric is calculated by taking the $Log10$ of the NA binder yield ratio, with the numerator as the yield of NA binder after entering "NA" and the denominator as the yield after entering "Pro".
@@
To further validate the effectiveness of the generated molecules, we selected two DNA targets (1C9Z[51], 1EEL[52]) and two RNA targets (2L94[53], 6XB7[54]) for docking experiments using NLDock[55], which is specially developed for simulate the interaction between nucleic acids and ligands. The docking process was configured in the local and flexible mode. We used a virtual screening dataset consisting of 6,400 molecules generated from a well-trained generator (Figure 2B). The results showed that for three out of the four targets (1EEL, 2L94, and 6XB7), many of our generated molecules exhibited stronger binding affinities than the original binders in their native conformations. Besides, compared with the original molecule, the structures are very diverse. However, for 1C9Z, due to its unique pocket shape, accurate molecular conformation generation played a crucial role in docking performance. When using only the SMILES of the native small molecule as input without specifying its experimental binding conformation, some of our generated molecules displayed superior affinity performance than the native ligand of this nucleic acid target.
It is worth noting that, compared to real molecules, the model struggles to generate highly symmetrical structures, such as those found in 1C9Z, 1EEL, and 2L94. Additionally, certain targets, such as 1PBR [56] and 1QD3 [57], have small-molecule ligands that entirely lack aromatic rings. However, in our experiments, we did not observe any generated samples without aromatic rings or double bonds. This limitation may stem from the scarcity of such molecules in the training dataset, making it difficult for the model to learn their characteristics. The proportion of fused aromatic rings tends to decrease because the SMILES notation for such structures is more complex. For example, we find molecules with fused aromatic rings targeting 1C9Z most in the first 25 epochs.
@@
GANs are prone to mode collapse and overfitting during training, often resulting in repetitive sequences such as "c1ccccc1" and "NCCCNCCCN." Unlike traditional SeqGAN and ORGAN, we applied a min-max transformation to map the rewards from both $D_{\psi}$ and $D_{n}$ to the range [0,1], preventing excessive rewards for certain sequences or bias toward a single discriminator (Figure 3A). This adjustment stabilized the training process, making the length fluctuate less.
@@
RL is highly sensitive to random seeds, with different seeds potentially leading to vastly different results[58, 59]. By comparing samples generated from different random seeds (Figure 3C,D), we observe that while all molecules are classified as nucleic acid binders, their specific structures vary significantly. To achieve greater molecular diversity, we strongly recommend conducting experiments with multiple random seeds to obtain a broader range of generated molecules.
@@ 4.2 CNS Drug Generation
Despite the pressing need, the number of approved CNS drugs remains limited. One major obstacle in CNS drug development is the blood-brain barrier (BBB), a selective membrane that restricts the entry of many compounds into the brain, complicating the delivery of therapeutic agents. Recent advancements in artificial intelligence and machine learning offer promising avenues to overcome these hurdles by enabling the design of novel compounds with optimized properties for CNS activity, but the few samples are still a problem[60]. We demonstrate the potential of our small sample based generative model to expand the library of candidate molecules for CNS drugs.
Oversampling Enable ADSeqGAN to Generate CNS Drugs
Figure 4: Changes in the yield of CNS drugs with training under different sampling methods
@@
To address this issue, we employed an oversampling technique during mixed training by tripling the sampling frequency of CNS drug data. This approach significantly improved the yield of CNS drugs, reaching over 5%, which is substantially higher than training without oversampling. However, further increasing the yield remains challenging due to two main reasons: (1) molecules of other classes also receive rewards, thus, the model tends to learn majority-class features to maximize local reward when sample distribution is imbalanced; (2) the limited number of CNS drug samples results in a less effective classifier during pretraining, as illustrated in Figure S4.
@@
Overall, ADSeqGAN combined with oversampling provides a novel strategy for generating molecules from small-sample datasets, demonstrating its potential for addressing data scarcity in drug discovery.
@@ 5 Conclusion
In this study, we proposed ADSeqGAN, a novel sequence-based GAN framework incorporating auxiliary discriminators for small-sample molecular generation. By integrating a pretrained classifier as an additional discriminator, ADSeqGAN improves the generation of specific molecular classes, i.e., nucleic acid binders and CNS drugs. Our results demonstrate that ADSeqGAN outperforms traditional non-pretrained generative models in terms of molecular validity, diversity, and class specificity.
Through a combination of MLE pretraining generator, Wasserstein loss, and data augmentation techniques such as oversampling, ADSeqGAN effectively addresses mode collapse and enhances the generative process. The model exhibits strong label responsiveness, successfully differentiating nucleic acid binders from protein binders and generating CNS drugs at a significantly higher rate than standard approaches. Furthermore, docking experiments confirm the ability of ADSeqGAN in generating high-quality molecules, thus it has a great application in enlarging compound library for nucleic acid drug discovery.
@@
Appendix
Appendix A Supplementary Information
This is appendix A.
Figure S1: Evaluation of random forest classifier of NA and Pro datasets
Figure S2: Principal component analysis of the characteristic molecular fingerprint fragments of NA dataset and generated NA samples with original Pro dataset
Figure S3: The impact of random number on the training process.
Figure S4: Evaluation of random forest classifier of CNS and non-CNS datasets

### v2

@@ 4.1 Nucleic Acid Binder Generation
To distinguish nucleic acid binders from protein-targeting small molecules, we computed molecular descriptors for a dataset of compounds and evaluated their discriminatory power using logistic regression. Following an automatical regression-based analysis, we manually selected 18 descriptors with distinct physicochemical properties that showed relevance in distinguishing these two classes from the top 50. A random forest classifier trained on these selected descriptors achieved an AUC of 0.91 (Figure S1), indicating strong classification performance.
Furthermore, we have integrated MACCS fingerprint56 based classifer into the pipeline, providing user-friendly choices and it shows slightly better performance than RDKit descriptors in Figure S5. MACCS is also more convenient, requiring no manual feature selection. We also tested deep learning–based descriptors (CDDD) 57, which can be directly extended to other chemical spaces, as shown in Figure S5. While CDDD provides the advantage of avoiding manual feature selection and offers broader transferability, it also introduces practical limitations, including compatibility issues, longer loading times for pretrained parameters, and increased computational cost. Therefore, for the current work, we retained RDKit/OpenBabel descriptors in the workflow in ADSeqGAN.
@@
Figure 2: ADSeqGAN training results on the nucleic acid and protein dataset. $A.$ The yield of nucleic acid binders using ”NA” as the input label changes with the increase of epoch. Yield is calculated by $unique\_ratio\times verified\_ratio\times NA\_ratio$. $B.$ NLDock docking results of generated molecules with various nucleic acid targets. 1C9Z and 1EEL are DNA targets, while 2L94 and 6XB7 are RNA targets. The grey balls are original binders with native conformations, and the red ones are generated samples. $C.$ Plot of label responsiveness with the increase of epoch. The responsiveness metric is calculated by taking the $Log10$ of the NA binder yield ratio, with the numerator as the yield of NA binder after entering ”NA” and the denominator as the yield after entering ”Pro”.
@@
To further validate the effectiveness of the generated molecules, we selected four DNA targets (1C9Z58, 1EEL59, 3U0560, 6AST61) and four RNA targets (1UUD62, 2L9463, 2LWK64, 6XB765) for docking experiments using NLDock45, which is specially developed for simulate the interaction between nucleic acids and ligands. The docking process was configured in the local and flexible mode. We used virtual screening datasets, with each consisting of 6,400 molecules generated from a well-trained generator and then chose 2 hits for each target as cases (Figure 2B, S6A and Table S1, S2). The results showed that for seven out of the eight targets, many of our generated molecules exhibited stronger binding affinities than the original binders in their native conformations. In addition, compared with the original molecule, the structures are very diverse. For 1C9Z, its unique pocket shape made it hard for generated molecule to perform better than ground truth. However, using only the native ligand’s SMILES (without experimental binding conformation), we successfully generated molecules—despite the difficulty of matching the pocket shape—that even showed higher affinity than the native ligand. AlphaFold3 was further used to confirm binding poses and observe local interactions. The coincidence of conformations predicted by AlphaFold3 and NLDock further enhances the reliability of the prediction results.
To further understand structural insight of generated molecules, we expanded our analysis to include scaffold distributions and functional group diversity as shown in Figure S7. Nucleic acid binders tend to contain a higher proportion of nitrogen and oxygen atoms and more fused aromatic ring systems, while protein binders more frequently feature aliphatic ring systems. Importantly, the molecules generated by ADSeqGAN captured these preferences, particularly the enrichment of aromatic rings and nitrogen atoms in nucleic acid binders.
It is worth noting that, compared to real molecules, the model struggles to generate highly symmetrical structures, such as those found in 1C9Z, 1EEL, and 2L94. Additionally, certain targets, such as 1PBR 66 and 1QD3 67, have small-molecule ligands that entirely lack aromatic rings. However, in our experiments, we did not observe any generated samples without aromatic rings or double bonds. This limitation may stem from the scarcity of such molecules in the training dataset, making it difficult for the model to learn their characteristics. The proportion of fused aromatic rings tends to decrease because the SMILES notation for such structures is more complex and difficult for GAN and RL to learn. Actually, we find molecules with fused aromatic rings targeting 1C9Z most in the first 25 epochs. Only certain random seeds (e.g., 666 and 888 after 40 reinforcement learning steps) yielded models capable of producing fused-ring structures. This reflects an inherent limitation of combining GANs with reinforcement learning, which may converge to local optima and struggle to capture highly complex structural features.
@@
GANs are prone to mode collapse and overfitting during training, often resulting in repetitive sequences such as ”c1ccccc1” and ”NCCCNCCCN.” Unlike traditional SeqGAN and ORGAN, we applied a min-max transformation to map the rewards from both $D_{\psi}$ and $D_{n}$ to the range [0,1], preventing excessive rewards for certain sequences or bias toward a single discriminator (Figure 3A). This adjustment stabilized the training process, making the length fluctuate less.
@@
RL is highly sensitive to random seeds, with different seeds potentially leading to vastly different results68, 69. By comparing samples generated from different random seeds (Figure 3C,D), we observe that while all molecules are classified as nucleic acid binders, their specific structures vary significantly. To achieve greater molecular diversity, we strongly recommend conducting experiments with multiple random seeds to obtain a broader range of generated molecules. As shown in Figure S7, seed 1111 produced scaffolds with few oxygen atoms and occasional macrocycles and seed 888 generated more linear structures, while seed 1234 yielded highly branched scaffolds. These findings are consistent with the structural diversity patterns shown in Figure 3D.
@@ 4.2 CNS Drug Generation
Figure 4: ADSeqGAN training results on the CNS and non-CNS drug dataset. $A.$ Influence of sampling strategies on CNS drug yield. $B.$ CNS_MPO statistic results of original dataset and four few-shot moleculr generation models. $C.$ Evaluation of metrics on few-shot generative algorithms to generate CNS drugs. $\Uparrow$ means the larger the better. $D.$ MOE docking result of generated molecules with various CNS targets. Green sticks show original molecules, while red sticks show generated samples. 4MM5 is LeuBAT (delta13 mutant) in complex with sertraline. 6AWP is ts3 human serotonin transporter complexed with fluvoxamine.
Despite the pressing need, the number of approved CNS drugs remains limited. One major obstacle in CNS drug development is the blood-brain barrier (BBB), a selective membrane that restricts the entry of many compounds into the brain, complicating the delivery of therapeutic agents. Recent advancements in artificial intelligence and machine learning offer promising avenues to overcome these hurdles by enabling the design of novel compounds with optimized properties for CNS activity, but the few samples are still a problem70. CNSMolGen71 is developed for CNS drug generation using a pretrained dataset based on either the central nervous system multiparameter optimization (CNS_MPO) score 72, 73 or high bioactivatity, followed by a fine-tuning module focusing on specific drug class. Our model goal is to build on the foundation of drugs that have entered clinical trials and have already been approved. We demonstrate the potential of our small sample based generative model to expand the library of candidate molecules for CNS drugs. To estimate the quality of generated samples, score functions designed for synthesizability (SA),and CNS_MPO are used. It should be noted that the calculation of CNS_MPO depends on basic pKa. To facilitate large-scale computing, we used MolGpKa74 based on graph neural networks for batch pKa prediction.
To evaluate performance under extremely few-shot conditions for CNS drug generation, we conducted comparative experiments using ADSeqGAN, CLM, and MolGen, all of which support molecular generation in low-data regimes. The training configurations for all models were kept identical to those used in the nucleic acid binder generation task. For MolGen, 50 new molecules were generated per input molecule in the CNS-targeting task.
Pretrain Auxiliary Descriminators
Overall, we adopted the same pipeline as previously used for constructing the nucleic acid-based auxiliary discriminator. Specifically, we first selected 12 molecular descriptors with strong physicochemical discriminability via logistic regression, and then trained a random forest classifier. The resulting classifier achieved an AUC of 0.81 (Figure S4), which is notably lower than the 0.91 obtained for the nucleic acid binder task. This performance gap can be attributed to two factors: (1) the physicochemical properties of CNS and non-CNS drugs are more similar than those between nucleic acid binders and approved drugs; and (2) the number of CNS drug samples is relatively limited. Therefore, directly using this classifier as a discriminator poses certain limitations. To enhance the model’s ability to capture CNS-specific characteristics, we adjusted the reward threshold such that a molecule is assigned a reward value of 1 if the classifier predicts a CNS probability greater than 40%.
Performance of ADSeqGAN to Generate CNS Drugs
@@
To address this issue, we employed an oversampling technique during mixed training by tripling the sampling frequency of CNS drug data. This approach significantly improved the yield of CNS drugs, reaching over 5%, which is substantially higher than training without oversampling (Figure 4A). However, further increasing the yield remains challenging due to two main reasons: (1) molecules of other classes also receive rewards, thus, the model tends to learn majority-class features to maximize local reward when sample distribution is imbalanced; (2) the limited number of CNS drug samples results in a less effective classifier during pretraining, as illustrated in Figure S4.
To further evaluate the quality of the generated compounds, we employed CNS_MPO as a metric (Figure 4B). The average CNS_MPO score of CNS drugs in the training set was 4.31. Molecules generated by ADSeqGAN without oversampling, as well as those produced by CLM and MolGEN, exhibited CNS_MPO distributions comparable to that of the training set. However, we observed a decrease in CNS_MPO scores when ADSeqGAN was trained with 3× oversampling. This decline may be attributed to the tendency of GANs to converge to a local optimum. Nonetheless, the decrease is acceptable, as ADSeqGAN with oversampling demonstrated a notably higher degree of molecular novelty compared to the other two few-shot models (Figure 4C). Furthermore, when the generated molecules were classified using our pre-trained random forest auxiliary discriminator, ADSeqGAN yielded significantly more molecules that matched the classifier’s criteria than the other models. It is important to note that the CNS_MPO score and the random forest classifier serve different purposes: CNS_MPO is designed to quantify the overall drug-likeness of CNS compounds, whereas the random forest model is trained to distinguish between CNS and non-CNS drugs. This difference in objectives may lead to a separation of shared features between CNS and non-CNS drugs that are otherwise embedded in the CNS_MPO metric, thereby contributing to the observed reduction in CNS_MPO scores.
To further validate the effectiveness of the generated molecules, we conducted molecular docking-based virtual screening against two CNS drug targets: 4MM5 75 and 6AWP 76. These experiments were designed to evaluate the binding ability of the generated compounds. A virtual screening dataset comprising 6,400 samples was constructed using a generator trained under optimal performance conditions. After docking, we chose top hits to do case studies (Figure 4D, S6B and Table S1, S2). The 4MM5 structure represents the complex of the antidepressant sertraline with LeuBAT. Among the generated compounds, we identified molecules with structures highly similar to sertraline. Docking results showed that the fluorophenyl group in the generated sample aligned closely with the dichlorophenyl moiety in sertraline, occupying nearly the same position within the binding pocket. Moreover, the secondary amine group of the sample molecule spatially overlapped with the amine nitrogen atom of sertraline. Notably, the docking score of the generated sample was slightly better than that of sertraline, suggesting comparable or enhanced binding affinity. We further used SiteAF377 to study binding poses and local interaction. In the 4MM5 system, ground-truth hydrogen bond distances were 4.2 Å and 3.4 Å, while our generated molecules exhibited 2.8 Å and 3.2 Å, showing superior binding ability. Similarly, 6AWP is the complex of the selective serotonin reuptake inhibitor (SSRI) fluvoxamine with the ts3 human serotonin transporter. Both the sample molecule and fluvoxamine featured a branched, tripodal structure. Docking analysis indicated that the amine group in the generated molecule occupied nearly the same spatial position as that in fluvoxamine and formed hydrogen bonds with the protein. In addition, the methoxy and aromatic side chains of both molecules exhibited similar orientations. Despite having comparable molecular sizes, the generated compound exhibited a higher docking score than fluvoxamine. Similar to 4MM5, the terminal amino groups of the molecules in 6AWP also capture similar hydrogen bond interactions. These findings support the capability of the ADSeqGAN model to effectively expand the chemical space of CNS-active molecules.
@@
Overall, both ADSeqGAN itself and combined with oversampling provide a novel strategy for generating molecules from small-sample datasets, demonstrating its potential for addressing data scarcity in drug discovery.
4.3 Expand ADSeqGAN Framework to CB1 ligand generation
To illustrate the versatility of ADSeqGAN, we showed how well ADSeqGAN perform in generating novel molecular binders for a specific drug target, here it is CB1. We have collected about three thousand CB1R ligands which have measured inhibition constant ki values. We first constructed a classification model using the MACCS fingerprint as the descriptors. The molecular dataset is roughly balanced in terms of the numbers of the active and inactive compounds if we applied 1 µM as the threshold. For the generated molecules, we applied a set of drug likeness filters including the QED score and Lipinski’s Rule of 5 to filter out those non-druglike molecules. We have reported that the Glide docking score is a poor predictor for CB1R, thus, applied a target-specific scoring function, ligand-residue interaction profile-scoring function (LRIP-SF),78 to prioritize the designed molecules. As detailed in the Supporting Text, the LRIP-SF achieved an encouraging scoring power with the root-mean-square error of 1.27 kcal/mol and ranking power with correlation coefficient of 0.64. We then applied the established LRIP-SF scoring function to predict the binding affinities of the designed druglike molecules. If we applied a threshold of -8.1854 kcal/mol, which corresponds to 1 µM, to determine if a compound is active or inactive, 32.8% of designed molecules belong to the active group. This is an encouraging performance as the hit rate surpasses most CB1R-focused compound libraries not to mention the general-purpose screening libraries. As a conclusion, ADSeqGAN is able to generate novel and druglike molecules for a specific drug target.
@@ 5 Conclusion
In this study, we proposed ADSeqGAN, a novel sequence-based GAN framework incorporating auxiliary discriminators for small-sample molecular generation. By integrating a pretrained classifier as an additional discriminator, ADSeqGAN improves the generation of specific molecular classes, i.e., nucleic acid binders, CNS drugs and CB1 ligands. Our results demonstrate that ADSeqGAN outperforms traditional non-pretrained generative models in terms of molecular validity, diversity, and class specificity.
Through a combination of MLE pretraining generator, Wasserstein loss, and data augmentation techniques such as oversampling, ADSeqGAN effectively addresses mode collapse and enhances the generative process. The model exhibits strong label responsiveness, successfully differentiating nucleic acid binders from protein binders and generating CNS drugs at a significantly higher rate than standard approaches. Furthermore, docking experiments confirm the ability of ADSeqGAN in generating high-quality molecules, thus it has a great application in enlarging compound library for nucleic acid and CNS drug discovery.
@@
6 Supporting Information
Additional experimental details on nucleic acid binders, central nervous system drugs, and CB1R ligands; evaluation of random forest classifiers built on physical chemical descriptors, MACCS and CDDD; PCA analysis; docking and binding pose comparisons; scaffold and functional group analysis; case study on CB1R with LRIP-SF; computational details of ADSeqGAN training and generation. (PDF)
7 Acknowledgment
This work was supported by funds from the National Institutes of Health (R01GM147673, R01GM149705) and the National Science Foundation (1955260). The authors would like to thank the computing resources provided by the Center for Research Computing (facility RRID: SCR_022735) at the University of Pittsburgh (NSF award number OAC-2117681), and the Pittsburgh Supercomputer Center (grant number BIO210185)
8 Data and Software Availability
All the training datasets, code and parameters are available online at GitHub: https://github.com/ClickFF/ADSeqGAN and https://github.com/HaCTang/ADSeqGAN.
9 AUTHOR INFORMATION
9.1 Corresponding Author
Junmei Wang - Department of Pharmaceutical Sciences and Computational Chemical Genomics Screening Center, School of Pharmacy, University of Pittsburgh, Pittsburgh, PA 15261, USA. Email: juw79@pitt.edu
9.2 Other Authors
Haocheng Tang - Department of Pharmaceutical Sciences and Computational Chemical Genomics Screening Center, School of Pharmacy, University of Pittsburgh, Pittsburgh, PA 15261, USA. Email: hat170@pitt.edu
Jing Long - School of Software & Microelectronics, Peking University, Beijing, 100871, CN. Email: jing.long0926@gmail.com
Beihong Ji - Department of Pharmaceutical Sciences and Computational Chemical Genomics Screening Center, School of Pharmacy, University of Pittsburgh, Pittsburgh, PA 15261, USA. Email: bej22@pitt.edu
9.3 Author Contributions
J.W. and H.T. designed the research. H.T., L.J. and B.J. performed the research. H.T. analyzed the data and wrote the manuscript. J.W. provided computing resources. All authors reviewed the manuscript.
9.4 Conflict of Interest
The authors declare no competing financial interests.